# Fast Hip Joint Moment Estimation with A General Moment Feature Generation Method

### v1

@@ Abstract
The hip joint moment during walking is a crucial basis for hip exoskeleton control. Compared to generating assistive torque profiles based on gait estimation, estimating hip joint moment directly using hip joint angles offers advantages such as simplified sensing and adaptability to variable walking speeds. Existing methods that directly estimate moment from hip joint angles are mainly used for offline biomechanical estimation. However, they suffer from long computation time and lack of personalization, rendering them unsuitable for personalized control of hip exoskeletons. To address these challenges, this paper proposes a fast hip joint moment estimation method based on generalized moment features (GMF). The method first employs a GMF generator to learn a feature representation of joint moment, namely the proposed GMF, which is independent of individual differences. Subsequently, a GRU-based neural network with fast computational performance is trained to learn the mapping from the joint kinematics to the GMF. Finally, the predicted GMF is decoded into the joint moment with a GMF decoder. The joint estimation model is trained and tested on a dataset comprising 20 subjects under 28 walking speed conditions. Results show that the proposed method achieves a root mean square error of 0.1180 ± 0.0021 Nm/kg for subjects in test dataset, and the computation time per estimation using the employed GRU-based estimator is 1.3420 ± 0.0031 ms, significantly faster than mainstream neural network architectures, while maintaining comparable network accuracy. These promising results demonstrate that the proposed method enhances the accuracy and computational speed of joint moment estimation neural networks, with potential for guiding exoskeleton control.
Index Terms:
@@
Joint moment, kinetics estimation, lower-limb exoskeleton, deep learning
I Introduction
@@
Lower-limb exoskeletons are capable of assisting in human locomotion, thus reducing joints moments, metabolic expenditure, and mechanical loading of joints. This allows it to be applied to assist weight-bearing workers, elderly individuals in need of walking assistance, and patients with arthritis [1, 2, 3, 4, 5]. During walking, the hip joint contributes 40-50% of the total positive joint power of the lower limb joints [6], which means that hip-assisted walking exoskeletons can bring substantial benefits. Furthermore, compared with other lower limb exoskeletons, hip exoskeletons have smaller metabolic penalties caused by distal-borne mass [7]. Therefore, hip exoskeletons are considered an important choice for walking assistance.
@@
The moment of the hip joint is an important basis for hip exoskeleton assistance. Since calculating joint moment with motion capture systems or force platforms is not feasible for real-world using of hip exoskeleton, it is necessary to utilize other sensors fixed on the exoskeleton system. Recent researchers have focused on determining the assistive torque of the exoskeleton based on the human motion information measured by these sensors, such as encoders on the motors or inertial measurement units (IMUs).
@@
Currently, there are two main methods to generate the assistive torque of hip exoskeletons based on hip joint motion data: 1) determining output torque based on assistive torque curve generation and gait phase estimation, and 2) calculating assistive torque according to instantaneous joint moment estimates. The former method involves the following steps: a) estimating the gait phase and the assistive torque profile for the entire gait cycle based on the user’s historical walking data, and b) generating the assistive torque based on the gait phase and the assistive torque profile [8, 9, 10]. However, this method may exhibit delayed assistive torque when dealing with variable walking gaits, which can hinder the user’s movements. The other method generates assistive torque directly through a computational model by measuring the kinematic information of the human body [11, 12, 13]. Compared to the first method, the second method calculates assistive torque independent of gait periodicity, allowing it to generate appropriate walking assistance in real-time even with changing gaits. Therefore, joint moment estimates-based assistive torque calculation has the potential to become an important computational method for future exoskeleton control.
@@
The estimation of hip biomechanical joint moment is most crucial in the instantaneous joint moment estimates-based method. However, it remains a significant challenge to estimate hip joint moment based on measurable data from the human body. Although some researchers have obtained joint moments by establishing approximate dynamic models [14, 15], these methods rely on complex sensor equipment such as plantar pressure sensors, which may decrease the stability of the exoskeleton system and reduce user comfort.
@@
As data-driven methods have shown remarkable effectiveness in biological measurement and exoskeleton control fields [16, 17, 18, 19], researchers have started focused on measured motion information (such as hip joint position changes) by simple and lightweight sensors during human walking for the estimation of joint moment. Wang used the XGBoost method to achieve the prediction of the knee joint adduction moment in an experiment involving 106 participants by using only the information from two IMU sensors [20]. However, traditional machine learning methods have limited performance in improving prediction accuracy of joint moment estimation, and researchers are turning to neural networks to further enhance the prediction accuracy of joint moment. Molinaro trained a TCN network that utilized expanded causal convolutional layers to encode temporal information and learn data features from input, achieving hip joint moment prediction based on data of IMU and angle of hip joint [21, 22]. Furthermore, Hossain designed an Ensemble network that improved the performance of human dynamics prediction using deep neural network layers with different structure types. Through fusion modules and bagging techniques, the model achieved a more accurate prediction of lower-limb joint moment [23]. However, the aforementioned joint moment estimation methods, which are mainly used for offline biomechanical joint moment estimation, suffer from the problem of long computation time and lack of personalization, which leads to their inapplicability to new users wearing hip exoskeletons.
To solve the above problems, we propose a fast joint moment estimation method based on general moment feature (GMF) generation. In the training phase, this method utilizes the user’s body parameters and hip joint moment to generate GMF using a feature generator. This GMF is then used to train a GRU (gated recurrent unit) neural network-based estimator, which can quickly predict the GMF. Additionally, a decoder is trained to decode the GMF into biomechanical joint moment. During the predicting phase, the trained estimator is used to rapidly estimate the GMF, which is subsequently decoded by the decoder to obtain the predicted joint moment.
Figure 1: The proposed method for fast hip joint moment estimation. During the training process, kinematic information ($\boldsymbol{X}$) is utilized as input for the estimator, while body parameters ($\boldsymbol{Q}$) and real joint moments ($M$) serve as inputs for the generator. The decoder takes body parameters and the GMF generated by the generator (represented by ”*”) as inputs. The trained model is then employed to estimate hip joint moments ($\hat{M}$) during walking for new users. It is worth noting that the generator does not exist in the prediction process, and the input to the moment decoder comes from estimator, which is different compared to the training process.
The main contributions of this study are:
@@
    Proposing a GMF generation method based on body parameters, which is conducive to obtaining a more individual-independent generalized representation of biomechanical joint moments and improves the performance of the joint moment estimation model.
@@
    Developing an estimator based on GRU, which addresses the real-time requirements of lower-limb exoskeleton application by taking advantage of the characteristics of recurrent neural networks in time series data prediction.
  • 3)
    Demonstrating the effectiveness of the proposed method through experiments with a dataset, which shows that it has higher accuracy, faster computation time than existing methods, and strong generalization capability.
@@
The structure of this paper will be described as follows. Section II describes the proposed fast joint moment estimation method based on GMF generation. The experiments and results are presented in sections III and IV, discussed in section V. The conclusion of this paper is written in section VI.
II Method
@@
This section provides a detailed explanation of the proposed fast joint moment estimation method based on GMF generation. We will firstly describe the basic model of using neural networks to estimate biomechanical joint moment. Then we introduce the concept of GMF, addressing the issues that arise from the general form. Based on this, we propose a network architecture for generating GMF as well as a decoder for it (Fig.1 - Model training) and introduce a fast estimator using GRU neural networks. Finally, we present the training method for the neural network model and the prediction process of estimating joint moment (Fig.1 - Prediction).
II-A The Basic Model of Joint Moment Estimation
@@
This study aims to estimate the hip joint moment based on the joint kinematics during walking utilizing neural networks. As mentioned in [24], the interpretability of a model is crucial for generating models with superior accuracy and robustness. Therefore, we first assume two essential attributes that a model should possess to estimate joint moment: 1) incorporating the relationship between kinematics and dynamics, and 2) ensuring the continuity of motion.
@@
The rationality of the second attribute is supported by previous researches [25, 26, 27], which indicates that using an appropriate window to capture the input time series can improve the prediction accuracy of the joint estimation model. Regarding the first attribute, to determine the model inputs from candidate kinematic information, we conducted a series of preliminary experiments. Finally, we selected six-dimensional data comprising the angles, angular velocities, and angular accelerations of the bilateral hip joints as the inputs for each time step of the estimator. The results of these experiments demonstrated that as the dimensionality of the inputs increased, the predictive accuracy of the joint moment estimation model gradually improved. Therefore, the general inputs of the model can be described as follows: $\boldsymbol{X}=[\boldsymbol{x}_{t},\boldsymbol{x}_{t-1},...,\boldsymbol{x}_{t-\Delta T}]\in R^{\Delta T\times D}$, where $\Delta T$ represents the window size and $D$ denotes the dimensionality of the joint kinematics data ($\Delta T=100$, $D=6$). A basic neural network model for estimating joint moment is described as $F$. For the estimated joint moment $\widehat{M}$ at time $t$, the process of estimation can be expressed as:
$\widehat{M}=\boldsymbol{F}(\boldsymbol{X}).$ (1)
II-B Generation of GMF
@@
The conventional form of joint moment estimation networks does not incorporate personalized information of the subjects. Therefore, to address the issue of decreased prediction accuracy when directly estimating joint moment using joint kinematics in a multi-user environment, we propose an estimation method based on GMF generation. Firstly, we introduce the concept of GMF, which is an individual-independent representation from the joint moment that combines with subject’s body parameters (mass and height are used in this paper). Next, a generator is proposed to generate this feature. Finally, we will explain the decoding process that decodes generated GMF into joint moment.
II-B1 GMF
Here, we first directly introduce the application form of GMF for joint moment estimation. According to the expression in (1), the joint moment estimation method based on GMF can be represented as follows:
$\widehat{GMF}=\boldsymbol{E}(\boldsymbol{X})$ (2)
$\widehat{M}=\boldsymbol{R}(\widehat{GMF}),$ (3)
@@ II-B1 GMF
where $\boldsymbol{E}$ denotes the GMF estimator and $\boldsymbol{R}$ denotes the GMF decoder, satisfying $\boldsymbol{R}(\boldsymbol{E}(\cdot))\in\boldsymbol{F}(\cdot)$. Using the estimator to estimate the proposed GMF instead of directly estimating the joint moment, we can reduce the decrease in prediction accuracy caused by individual differences, improving the generalization performance of the joint estimation model across multiple users.
@@
Based on the aforementioned requirements, we design a method for generating GMF and apply it to the training of the joint moment estimation model. Fig.1 illustrates the process of generating GMF. The joint moment estimation method based on GMF leverages both the joint moment and body parameters. It utilizes an adaptive neural network to extract GMF that is independent of individuals. This feature is then used to train the estimator. Additionally, a decoder is trained to decode the estimated GMF into joint moment. It is worth noting that, unlike directly inputting body parameters into the network, this method expresses the role of body parameters in the network model more specifically, which means it can achieve the optimal mapping between hip joint kinematics and joint moment.
II-B2 GMF Generator
@@
A feature generator $\boldsymbol{G}$ is used to generate the GMF. More specifically, this generator utilizes a neural network to combine the joint moment of subjects with their body parameters and finally generates the desired GMF.
@@
The main objective of GMF generation is to discover a better prediction target to achieve a more accurate estimation by the estimator $\boldsymbol{E}$ as described by (2). Therefore, unlike typical deep learning tasks, in this method, the labels learned by the estimator and the GMF generator are mutually the output values of each other during the training process. This structure guides the GMF generator to generate training labels, namely the proposed GMF, from the joint moment and body parameters of subjects, which can enhance the accuracy of the estimator.
@@
At the same time, it is notable that the GMF generated by the generator should have the capability to be decoded into joint moment. Using $\boldsymbol{\theta}_{\boldsymbol{G}}$ to represent the network parameters of the feature generator $\boldsymbol{G}$, $\boldsymbol{Q}$ and $M$ to represent the values of the subject’s body parameters and joint moment, respectively ($\boldsymbol{Q}=[m,h]$), the process of generating GMF can be expressed as follows:
$GMF=\boldsymbol{G}(\boldsymbol{Q},M;\boldsymbol{\theta}_{\boldsymbol{G}}).$ (4)
The optimal generator parameters $\boldsymbol{\theta}_{\boldsymbol{G}}^{*}$ in the above process can be mathematically represented as:
@@
$\boldsymbol{\theta}_{\boldsymbol{G}}^{*}=\mathop{\textbf{argmin}}\limits_{\boldsymbol{\theta}_{\boldsymbol{G}}}$ $\left\|\boldsymbol{E}(\boldsymbol{X};\boldsymbol{\theta}_{\boldsymbol{E}})-\boldsymbol{G}(\boldsymbol{Q},M;\boldsymbol{\theta}_{\boldsymbol{G}})\right\|_{2}$ (5)
$s.t.~{}\boldsymbol{G}(\boldsymbol{\theta}_{\boldsymbol{G}})\in\boldsymbol{S}(\boldsymbol{\theta}_{S}),$
@@
where $\boldsymbol{S}$ represents the function space in which $\boldsymbol{G}$ resides, ensuring the GMF generated through $\boldsymbol{G}$ can be decoded into joint moment.
@@
The network structure of training phase (as shown in Fig. 1) provides a specific illustration of the role of the GMF generator within the overall network structure. The inputs of the generator include joint moment and body parameters, and its output is the GMF. The GMF Generator serves two important functions: a) generating the optimal GMF to improve the prediction accuracy of the estimator, enhancing the model’s performance across different users; b) ensuring the decoder to accurately decode the generated GMF into joint moment.
To implement an model with the proposed method, as an example, we employ a fully connected network with five hidden layers to construct the GMF generator. Each hidden layer consists of 32 nodes, and LeakyReLU activation function is applied after each hidden layer to introduce non-linearity.
II-B3 GMF Decoder
The GMF decoder is used to decode the generated GMF into joint moment. We employ the same network structure as the GMF generator for constructing the GMF decoder. To improve the accuracy of the output, we include body parameters as part of its inputs. Using $\boldsymbol{R}$ to represent the moment decoder, $\boldsymbol{\theta}_{\boldsymbol{R}}$ as its network parameters, and $\widehat{M}$ as the decoded joint moment values, the process of joint moment decoding can be expressed as:
$\widehat{M}=\boldsymbol{R}(\boldsymbol{Q},GMF;\boldsymbol{\theta}_{\boldsymbol{R}}).$ (6)
The optimal parameters $\boldsymbol{\theta}_{\boldsymbol{R}}^{*}$ of the GMF decoder in the above process can be obtained from:
$\boldsymbol{\theta}_{\boldsymbol{R}}^{*}=\mathop{\textbf{argmin}}\limits_{\boldsymbol{\theta}_{\boldsymbol{R}}}\left\|\boldsymbol{R}(\boldsymbol{Q},GMF;\boldsymbol{\theta}_{\boldsymbol{R}})-M\right\|_{2}.$ (7)
II-C GMF Estimator
To ensure the efficiency of the model, we design the GMF estimator with more consideration to the computational efficiency of the network, using a GRU neural network.
Since the inputs ($\boldsymbol{X}$) and outputs ($\widehat{GMF}$) of the model in this study are both time series, joint moment estimation is essentially a time series regression problem. In such problems, recurrent neural networks have demonstrated high accuracy [28]. In the designed exoskeleton system, these networks can better leverage their ability to estimate joint moment in real-time. Based on the characteristic of memory information propagation, the network only needs to inherit the hidden state information from the previous time step and input the features of the current time step during real-time computation. Therefore, the inputs of the model can be redefined as: $\boldsymbol{X}_{t}=[\boldsymbol{x}_{t},\boldsymbol{H}_{t}]\in R^{2\times D}$ (ideally, $[\boldsymbol{x}_{t},\boldsymbol{H}_{t}]\equiv[\boldsymbol{x}_{t},\boldsymbol{x}_{t-1},...,\boldsymbol{x}_{t-\Delta T}]$). The relationships between each time step are explicitly expressed without the need for internal processing within the network, which significantly reduces the number of network parameters.

### v2

# Joint Moment Estimation for Hip Exoskeleton Control: A Generalized Moment Feature Generation Method

@@ Abstract
Hip joint moments during walking are the key foundation for hip exoskeleton assistance control. Most recent studies have shown estimating hip joint moments instantaneously offers a lot of advantages compared to generating assistive torque profiles based on gait estimation, such as simple sensor requirements and adaptability to variable walking speeds. However, existing joint moment estimation methods still suffer from a lack of personalization, leading to estimation accuracy degradation for new users. To address the challenges, this paper proposes a hip joint moment estimation method based on generalized moment features (GMF). A GMF generator is constructed to learn GMF of the joint moment which is invariant to individual variations while remaining decodable into joint moments through a dedicated decoder. Utilizing this well-featured representation, a GRU-based neural network is used to predict GMF with joint kinematics data, which can easily be acquired by hip exoskeleton encoders. The proposed estimation method achieves a root mean square error of 0.1180 ± 0.0021 Nm/kg under 28 walking speed conditions on a treadmill dataset, improved by 6.5% compared to the model without body parameter fusion, and by 8.3% for the conventional fusion model with body parameter. Furthermore, the proposed method was employed on a hip exoskeleton with only encoder sensors and achieved an average 20.5% metabolic reduction ($p<0.01$) for users compared to assist-off condition in level-ground walking.
keywords:
@@
lower-limb exoskeleton , hip joint moment estimation , deep learning
[…]
1 Introduction
@@
Lower-limb exoskeletons are capable of assisting in human locomotion, thus reducing joint moments, metabolic expenditure, and mechanical loading of joints. This allows it to be applied to assist weight-bearing workers, elderly individuals in need of walking assistance, and patients with arthritis [1, 2, 3, 4, 5]. During walking, the hip joint contributes 40-50% of the total positive joint power of the lower limb joints [6], which means that hip-assisted exoskeletons can bring substantial benefits. Furthermore, compared with other lower limb exoskeletons, hip exoskeletons have smaller metabolic penalties caused by distal-borne mass [7]. Therefore, hip exoskeletons are considered an important choice for walking assistance.
@@
The moment of the hip joint is a fundamental basis for hip exoskeleton assistance control. Since calculating joint moments with motion capture systems or force platforms is not feasible for the hip exoskeleton in the real world, it is necessary to utilize other sensors fixed on the exoskeleton system. Recent researchers have focused on determining the assistive torque of the exoskeleton based on the human motion information measured by these sensors, such as encoders on the motors or inertial measurement units (IMUs).
@@
Currently, there are two main methods to generate the assistive torque of hip exoskeletons based on hip joint motion data: determining output torque based on assistive torque curve generation and gait phase estimation, and calculating assistive torque according to instantaneous joint moment estimates. The former method involves the following steps: (a) estimating the gait phase and the assistive torque profile for the entire gait cycle based on the user’s historical walking data, and (b) generating the assistive torque based on the gait phase and the assistive torque profile [8, 9, 10, 11]. However, this method may exhibit delayed assistive torque when dealing with variable walking gaits, which can hinder the user’s movements. The other method generates assistive torque directly through a computational model by measuring the kinematic information of the human body [12, 13, 14]. Compared to the first method, the second method calculates assistive torque independent of gait periodicity, allowing it to generate appropriate walking assistance in real-time even with changing gaits. Therefore, calculating assistive torque with instantaneous joint moment estimation has the potential to become an important computational method for future exoskeleton control.
@@
Estimatign the hip biomechanical joint moment accurately is the key foundation to calculate assistive torque instantaneously, but it remains a significant challenge to obtain the estimates based on measurable data from the wearable sensors on the human body. Although some researchers have calculated joint moments by establishing approximate dynamic models [15, 16], these methods rely on complex sensor equipment such as ground reaction force sensors, which increases the system complexity, leading to the instability of the exoskeleton system. Besides, extra sensors can be uncomfortable for the users to wear and even hinder their movement.
@@
As data-driven methods have shown remarkable effectiveness in biological measurement and exoskeleton control fields [17, 18, 19, 20], researchers have started focusing on measured motion information (such as hip joint position changes) by simple and lightweight sensors during human walking for the estimation of joint moment. Wang used the XGBoost method to achieve the estimation of the knee joint adduction moment in an experiment involving 106 participants by using only the information from two IMU sensors [21]. However, traditional machine learning methods have limited performance in improving the estimation accuracy of joint moment estimation, and researchers are turning to neural networks to further enhance the estimation accuracy of joint moment. Molinaro trained a TCN network that utilized expanded causal convolutional layers to encode temporal information and learn data features from input, achieving hip joint moment estimation based on data of simulated IMU and angle of hip joint [22, 11]. Furthermore, Hossain designed an Ensemble network that utilized deep neural network layers with different structure types, improving the performance of human dynamics estimation. Through fusion modules and bagging techniques, the model achieved a more accurate estimation of lower-limb joint moment [23].
For the hip exoskeleton control, the joint moment estimation model should be able to adapt to new users so it can provide accurate joint moment estimation in real-time. However, the existing methods have not yet achieved this goal, for the reasons that the influence of human parameters is underestimated or difficult to obtain an estimation model that integrates individual characteristics. The lack of personalization might lead to poor applicability to new users, where the assistance torque generated by the exoskeleton might not be suitable for the user’s walking.
To address the above problems, we propose a joint moment estimation neural network based on generalized moment feature (GMF) generation. A feature generator was designed to obtain GMF, which is more generalized than the joint moment, facilitating the estimator to learn the relationship between the kinematics and dynamics during human walking better. Meanwhile, the GMF can be decoded with body parameters into the joint moment, achieving personalizing joint moment estimation. Furthermore, a GRU-based estimator was designed to reduce model computation time with minor accuracy loss under the proposed GMF neural network framework.
The main contributions of this study are as follows:
@@
    Proposing a novel joint moment estimation neural network based on user-specific body parameters, which is conducive to obtaining a superior individual-independent representation of biomechanical joint moments and achieving better estimation accuracy.
@@
    Developing the proposed method on a hip exoskeleton to validate its assistance gain for users as well as showing its potential application value in scenarios relying solely on encoder sensors.
@@
The structure of this paper will be described as follows. Section II describes the proposed joint moment estimation method based on the GMF generation. The experiments and results are presented in sections III and IV, discussed in section V. The conclusion of this paper is written in section VI.
2 Method
@@
This section provides a detailed explanation of the proposed joint moment estimation method based on GMF generation. We will first describe the conventional form of using neural networks to estimate biomechanical joint moments. Then we introduce the concept of GMF, addressing the issues that arise from the conventional form. Based on this, we propose a neural network framework for generating GMF and estimating it using a GRU module, as well as a decoder for decoding it into the joint moment (Fig.1 - Model training). Finally, we present the training strategy for the proposed neural network model and the deployment of estimating joint moment in the real world (Fig.1 - Model deployment).
Figure 1: Joint moment estimation for hip exoskeleton assistance based on GMF generation. During the training process, kinematics information ($\boldsymbol{X}$) is utilized as input for the estimator, while body parameters ($\boldsymbol{q}$) and hip joint moments ($M$) serve as inputs for the generator. The decoder takes body parameters and the GMF generated by the generator (represented by ”*”) as inputs. The estimator and decoder are then employed to estimate hip joint moments ($\hat{M}$) to provide assistance torque during walking for new users (the generator is not needed).
2.1 The Basic Model of Joint Moment Estimation
@@
This study aims to estimate the sagittal hip joint moments based on the joint kinematics during walking utilizing neural networks. As mentioned in [24], the interpretability of a model is crucial for generating models with superior accuracy and robustness. Therefore, two essential attributes are assumed that a neural network model should possess to estimate joint moment: (a) incorporating the relationship between kinematics and dynamics, and (b) ensuring the continuity of motion.
@@
The rationality of the second attribute is supported by previous researches [25, 26, 27], which indicates that using an appropriate window to capture the input time series can improve the estimation accuracy of the joint estimation model. Regarding the first attribute, to determine the model inputs from candidate kinematic information, we conducted a series of preliminary experiments. Data comprising the angles, angular velocities, angular accelerations, and their combinations of the bilateral hip joints were selected as the inputs for each time step of the estimator. The results of these experiments demonstrated that as the dimensionality of the inputs increased, the accuracy of the joint moment estimation model gradually improved. Therefore, the general inputs of the joint moment estimation model can be described as follows: $\boldsymbol{X}=[\boldsymbol{x}_{t},\boldsymbol{x}_{t-1},...,\boldsymbol{x}_{t-\Delta T}]\in R^{\Delta T\times D}$, where $\Delta T$ represents the window size and $D$ denotes the dimensionality of the joint kinematics data ($\Delta T=100$, $D=6$). A basic neural network model for estimating joint moments is described as $F$. For the estimated joint moment $\widehat{M}$ at time $t$, the process of estimation can be expressed as:
$\widehat{M}=F(\boldsymbol{X}).$ (1)
2.2 GMF-based Model
@@
The conventional form of joint moment estimation networks does not incorporate personalized information of the subjects as the individual parameters are not included in (1). Therefore, to address the issue of decreased accuracy when estimating joint moments using only joint kinematics in a multi-user environment, we propose an novel joint moement estimation method based on GMF generation.
2.2.1 Inspiration of GMF Generation
Firstly, we introduce the concept of GMF, which is an individual-independent representation from the joint moment that combines with the subject’s body parameters $\boldsymbol{q}$ (mass $m$ and height $h$ are used in this paper, which means $\boldsymbol{q}=[m,h]$). Similar to the expression in (1), the joint moment estimation method based on GMF can be represented as follows:
$GMF=G(\boldsymbol{q},M)$ (2)
$\widehat{GMF}=E(\boldsymbol{X})$ (3)
$\widehat{M}=D(\widehat{GMF}),$ (4)
where $M$ denotes the ground truth of joint moment, $G$ denotes the GMF generator, $E$ denotes the GMF estimator and $D$ denotes the GMF decoder, satisfying $D(E(\cdot))\in F(\cdot)$.
@@ 2.2.1 Inspiration of GMF Generation
Equations of (2), (3), and (4) show the core framework of GMF-based model to achieve a better estimation: estimating the GMF instead of directly estimating the joint moment, so that the decrease in estimation accuracy caused by individual differences can be reduced. In other words, GMF-based model improves the generalization performance of the joint moment estimation across multiple users.
@@
Based on the these inspiration, we design a method for generating GMF and apply it to the training of the joint moment estimation model. Fig. 1 illustrates the process of generating GMF. The joint moment estimation method based on GMF leverages both the joint moment and body parameters. It utilizes an adaptive neural network to extract GMF that is independent of individuals. This feature is then used to train the estimator. Additionally, a decoder is trained to decode the estimated GMF into the joint moment. It is worth noting that, compared with inputting body parameters into the network directly, this method expresses the role of body parameters in the network model more specifically, which means it can achieve better optimization from training.
2.2.2 GMF Generator
@@
A feature generator $G$ is used to generate the GMF. More specifically, this generator utilizes a neural network to combine the joint moment of subjects with their body parameters and finally generates the desired GMF.
@@
The main objective of GMF generation is to discover a better estimation target to achieve a more accurate estimation by the estimator $E$ as described by (3). Therefore, unlike typical deep learning tasks, in this method, the labels learned by the estimator and the GMF generator are mutually the output values of each other during the training process. This structure guides the GMF generator to generate training labels, namely the proposed GMF, from the joint moment and body parameters of subjects, which can enhance the accuracy of the estimator.
@@
Meanwhile, it is essential that the GMF generated by the generator should be able to be decoded into the joint moment. Using $\boldsymbol{\theta}_{G}$ to represent the network parameters of the feature generator $G$, the optimal generator parameters $\boldsymbol{\theta}_{G}^{*}$ in the above process can be mathematically represented as:
@@
$\boldsymbol{\theta}_{G}^{*}=\mathop{\textbf{argmin}}\limits_{\boldsymbol{\theta}_{G}}$ $\left\|E(\boldsymbol{X})-G(\boldsymbol{q},M;\boldsymbol{\theta}_{G})\right\|_{2}$ (5)
$s.t.~{}G(\boldsymbol{\theta}_{G})\in\boldsymbol{\mathcal{G}},$
@@
where $\boldsymbol{\mathcal{G}}$ represents the function space to which $G$ belongs, ensuring the GMF generated through $G$ can be decoded into joint moment.
@@
The network structure of the training phase (as shown in Fig. 1) provides a specific illustration of the role of the GMF generator within the overall network structure. The inputs of the generator include joint moment and body parameters, and its output is the GMF. The GMF Generator serves two important functions: (a) generating the optimal GMF to improve the estimation accuracy of the estimator, enhancing the model’s performance across different users; (b) ensuring the decoder can accurately decode the generated GMF into the joint moment. To implement a model with the proposed method, as an example, we employ a fully connected network with five hidden layers to construct the GMF generator. Each hidden layer consists of 32 nodes, and a LeakyReLU activation function is applied after each hidden layer to achieve non-linearity.
2.2.3 GMF Estimator